OpenStudy (annie96):

Which is the equation of the line passing through (0,4) and (8,0)? y=-1/2x+4 y=-1/2x+8 y=-2x+4 y=-2x+8 Combine like terms to simplify the expression: 7x+3y-2+6x-1- 2 y

OpenStudy (anonymous):

m=(y2-y1)/(x2-x1) y=m(x-x1)+y1

OpenStudy (anonymous):

There is 2 methods to solve this.... 1) substitute two poins in the eqution if it satisfies then that line is passing through the points... 2) find the slope of the points and then compare with the slopes of the line...

OpenStudy (anonymous):

We're not supposed to give explicit answers. Sorry, if that's what you wanted, Annie.

OpenStudy (annie96):